# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

### v2

@@ 4.1 Tasks
[…]
@@
We evaluate MOCOBO on two variations of the peptide design task: “template free" (TF) and “template constrained" (TC). For TF, we allow the optimizer to propose any sequence of amino acids. For TC, we add a constraint that any sequence proposed by the optimizer must have a minimum of $75$ percent sequence similarity to at least one of the $10$ template amino acid sequences in Table A.8. These $10$ templates were mined from extinct organisms and selected by Wan et al. [66]. The motivation of the template constrained task is to design peptides specifically likely to evade antibiotic resistance by producing “extinct-like” peptides that bacteria have not encountered in nature in thousands of years. For MOCOBO and all BO baselines (TuRBO, LOL-BO, ROBOT, and MORBO), we handle the optimization constraint by adapting techniques from SCBO [16]. For CluSO and random sampling, we handle the constraint with rejection sampling.
@@
Figure 3: (Left) In vitro results for the two best “template free" (TF1, TF2) and two best “template constrained" (TC1, TC2) runs of MOCOBO for the peptide design task. Columns are the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target pathogenic bacteria B1,,$\ldots$,, … , B11 listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive respectively. TF1 and TC1 correspond to the single runs of MOCOBO shown in Table 1 and Table A.1 respectively. Methods used to obtain in vitro MICs are provided in Section A.1. (Right) Ablation study comparing MOCOBO to optimization performance where a known “good” partitioning of the $T$ objectives into $K$ subsets is available in advance. We individually optimize $K$ solutions, one for each partition.
@@
In high dynamic range (HDR) images, some pixels (often associated with light sources) can dominate overall contrast, requiring adjustments to reveal detail in low-contrast areas, a problem known as tone mapping [51, Section 6]. Tone mapping algorithms involve various tunable parameters, resulting in a high-dimensional optimization problem of subjectively perceived quality. We seek a covering set of $K=4$ solutions to optimize a set of $T=7$ image aesthetic (IAA) and quality (IQA) assessment metrics from the pyiqa library [9] (see metrics listed in Table A.5). Our practical goal is that, while we do not know a priori which metric is best for a particular image, covering all metrics may result in at least one high quality image. We optimize over the $13$-dimensional parameter space of an established tone mapping pipeline to tone-map the “Stanford Memorial Church" [12] and “desk lamp" [7] benchmark images. See Section A.8.1 for details.
@@ 4.2 Optimization results
In Figure 2, we provide optimization results comparing MOCOBO to the baselines discussed above on all tasks. The results show that MOCOBO finds sets of $K$ solutions that achieve higher coverage scores across tasks. The “T Individually Optimized Solutions" baseline appears as a horizontal dotted line in all plots of Figure 2, representing the average coverage score of $T$ individually optimized solutions, serving as an approximation of the best possible performance without the constraint of a limited $K<T$ solution set. Results in Figure 2 demonstrate that the smaller set of $K<T$ solutions identified by MOCOBO nearly equals the performance of the complete set of $T$ individually optimized solutions. This result depends on using domain knowledge to choose $K$ large enough to achieve it. Results in Figure 2 show that matching the performance of $T$ optimized solutions is possible with some values of $K\ll T$.
@@
In Table 1, we provide the $K=4$ peptides found by one run of MOCOBO for the template free (TF) peptide design task. For each peptide, we provide the the APEX 1.1 model’s predicted MIC for each of the $11$ target bacteria. MIC values $\leq 16$ $\mu$ mol $L^{-1}$ are considered to be “highly active" against the target pathogenic bacteria [66]. We highlight in Table 1 the comparison between the second peptide with the other three peptides. B1-B7 are Gram negative (GN) bacteria, while B8-B11 are Gram positive (GP). Peptide 2 specialized to the GP bacteria (B8-B11 scores predicted highly active compared to the other peptides) at the expense of broad spectrum activity for the GN bacteria (B5, B6, B7 predicted inactive). By specializing its solutions to GN and GP bacteria separately, MOCOBO achieves low MIC across the target bacteria with only $K=4$ peptides. A similar table of results is also provided for the template constrained (TC) peptide design task (see Table A.1).
@@
In Figure 3 left, we provide in vitro results for the two best TF and two best TC runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. For each run, Figure 3 left provides the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target bacteria. Results demonstrate that solutions found by MOCOBO optimizing against the in-silico APEX 1.1 model achieve good coverage of the $11$ target pathogenic bacteria in vitro. In Figure A.1, we provide the full set of in vitro results for each these runs of MOCOBO, with MIC values obtained by each of the $K=4$ peptides found for each target bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
@@
In Table A.2, we provide results for the $K=3$ molecules found by one run of MOCOBO for the molecule design task. Each of the $6$ target elements is successfully present in one of the $K=3$ molecules in the best covering set. These three molecules therefore effectively cover the $T=6$ objectives, with each objective having a max score $>0.9$.
@@
In Figure 4 left, we depict the $4$ varied obstacle courses used for the $T=4$ variation of the rover task and $K=2$ covering trajectories found by a single run of MOCOBO. In Figure A.3 and A.4, we provide analogous figures for the $T=8$ and $T=12$ variations. In each example, the MOCOBO optimized set of trajectories covers all obstacle courses such that all obstacles are avoided.
@@
In Figure 4 right and Figure A.2, we provide the original hdr images and the images obtained by the $K=4$ solutions found by a single run of MOCOBO for the church and desk test images respectively. In both variations, three of the four solutions result in high quality images, while one results in a poor quality image (see middle right church image and rightmost desk image). The poor quality images were each generated by the one solution that MOCOBO used to cover metrics $5$ and $6$ (see metric ID numbers in Table A.5). This indicates that metrics $5$ and $6$ are poor indicators of true image quality in this setting. This result highlights a useful application of MOCOBO: By dedicating one of the $K$ solutions to maximizing the misleading metrics ($5$ and $6$), the other three solutions to can focus on the remaining subjectively better metrics.
@@ 4.3 Ablation study
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (1) reduces to optimizing the sum of objectives in each relevant subset independently. In Figure 3 right, we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization. See Section A.2.2 for an additional ablation study of MOCOBO when all objectives are highly conflicting.
@@ 5 Discussion
By bridging the gap between traditional multi-objective Bayesian optimization (BO) and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions. See Section A.4 for additional discussion and limitations of MOCOBO.
@@ A.2.1 Additional covering sets of solutions found by MOCOBO
In this section, we provide additional examples of covering sets of solutions found by MOCOBO for various tasks from Section 4.
[…]
In Figure A.1, we provide In vitro results for the two best “template free" and two best “template constrained" runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. Figure A.1 provides in vitro MICs for each of the $K=4$ peptides found by each of these runs of MOCOBO, for each of the $11$ target pathogenic bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
@@
Table A.1: The best set of $K=4$ peptide sequences found by one run of MOCOBO for the “template free" peptide design task described in Section 4.1. For each of the four sequences, we provide the MIC according to the APEX model for each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TC1" in Figure 3 left, and Figure A.1 for in vitro MICs for this set of $K=4$ peptides.
In Table A.1, we provide an example of a covering set of peptides found by MOCOBO for the “template constrained" variation of the peptide design task.
In Figure A.2, we provide the original hdr image (leftmost panel), and the $K=4$ images produced using the $K=4$ solutions found by a single run of MOCOBO for the church image variation of the image tone mapping task. An analogous result for the desk image variation of the image tone mapping task can be found in the main text in Figure 4 right. A notable apparent limitation of the existing IAA/IQA metrics we used for the image tone mapping tasks is that they all favored monochromatic images for both the church and desk images. As such, the tone-mapped images obtained appear less colorful than their hand-tuned counterparts reported in prior work.
In Table A.2, we provide an example of a covering set of $K=3$ molecules found by MOCOBO for the molecule design task. As mentioned in Section 4, these three molecules effectively cover the $T=6$ objectives, as evidenced by the presence of all $T=6$ target elements in one of the $K=3$ molecules designed by MOCOBO.
In Figure A.3 and Figure A.4, we provide examples of a covering set of trajectories found for the $T=8$ and $T=12$ variations of the rover task respectively. In each figure, a panel is shown for each of the $T$ obstacles courses, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. Each panel also shows the best among the $K$ trajectories found by a run of MOCOBO for navigating each obstacle course. An analogous plot for the $T=4$ variation of the rover optimization task is provided in the main text in Figure 4.
@@
Table A.2: The best set of $K=3$ molecules found by one run of MOCOBO for the Ranolazine MPO multiple element molecule design task described in Section 4.1. For each of the three molecules, we provide the objective value obtained for each of the $T=6$ objectives. The best/highest objective value is in bold in each column. Each objective aims to add a different target element to Ranolazine (F, Cl, Br, Se, S, and P). The $T=6$ target elements are in bold in the smiles string [69] representation of each molecule.
@@ A.2.2 Additional ablation study: $T$ highly conflicting objectives
MOCOBO is designed for cases where some, but not all, of the objectives are highly conflicting, and this is the case for all of the tasks considered in Section 4. The conflicting pairs of the objectives prevent a single solution from optimizing all objectives well. The fact that not all pairs of objectives are completely conflicting, is the reason why it is possible to cover all T objectives with a small set of K solutions. If all pairs of the T objectives are highly conflicting, there is by definition no possible set of K<T solutions such that all objectives are well optimized. In this case, it’s best to use T individually optimized solutions if your goal is to find at least one solution that well optimizes each objective. However, it is still interesting to consider the performance of MOCOBO in the setting of T objectives that are pairwise highly conflicting. To investigate this, we construct a new variation of the multi-objective rover task described in Section 4.1. We design T=3 obstacle courses such that it is impossible for the rover to take any single path that avoids all obstacles in any pair of the obstacle courses. This results in T=3 objectives that are pairwise highly conflicting. We then run 20 replications of MOCOBO on this problem with K=2, asking MOCOBO to design K=2 solutions that cover the T=3 pairwise highly conflicting objectives. For this task, MOCOBO got an average coverage score of $2.932$, while T individually optimized solutions got an average coverage score of $12.899$. It is unsurprising that T individually optimized solutions achieved better coverage here since all three objectives are pairwise highly conflicting. However, this result demonstrates that MOCOBO is able to design sets of K=2 solutions that still achieve fairly high coverage scores.
@@
Figure A.5 provides a diagram of the T=3 pairwise highly conflicting obstacle courses, and an example of one of the covering sets of K=2 trajectories found by a single run of MOCOBO. One optimized trajectory (shown in blue) is designed by MOCOBO such that it specializes to navigate the first obstacle course, navigating it without hitting any obstacles (Leftmost Panel). The other optimized trajectory (shown in magenta) is designed by MOCOBO such that it avoids all obstacles in the third obstacle course (Rightmost Panel), while also minimizing total amount of impact with the obstacles in the second obstacle course (Middle Panel). Since it is by-design impossible to avoid all obstacles in all T=3 obstacle courses with only K=2 solutions, Figure A.5 demonstrates that MOCOBO was able to successfully balance trade-offs among the objectives, designing a set of K=2 solutions that minimized the total amount of obstacle impact across the three obstacle courses.
A.2.3 Wall-clock runtime of Algorithm 1
In Table A.3, we provide an empirical evaluation of the average wall-clock runtime of Algorithm 1 with different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). We provide average run times with $N=20,200,2000,20000,200000,$ and $2000000$ for each combination of $K$ and $T$ that we ran MOCOBO on in Section 4. To improve the speed of Algorithm 1, in all experiments, we compute the inner for loop (lines 5-6 in Algorithm 1) in parallel. This involves simply computing the marginal coverage of adding all $N$ data points at once in parallel, and then updating $A$ with the point that achieved maximum marginal converge improvement. Parallel computation of the marginal coverage improvement of the $N$ points is straightforward to implement using PyTorch. In Table A.3, the “Runtime" column gives the average runtime of Algorithm 1 with this parallel computation of the of the inner for loop, and the “Runtime No Parallelization" column gives the average runtime without the parallel computation. Notice that as $N$ grows large, this parallel computation becomes essential to achieve the reasonably fast run times of Algorithm 1 needed to run MOCOBO. Average run times reported in Table A.3 are computed by running Algorithm 1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs.

### v3

@@ 4.1 Tasks
In the peptide design task, we explore amino acid sequences to minimize the MIC (minimum inhibitory concentration, measured in $\mu$ mol $\text{L}^{-1}$) for each of $T=11$ target drug resistant strains or otherwise challenging to kill bacteria (B1-B7 Gram negative, B8-B11 Gram positive). Table˜A.7 lists our target bacteria in this study. Briefly, MIC indicates the concentration of peptide needed to inhibit bacterial growth (see Kowalska-Krochmal and Dudek-Wicher 34). We evaluate MIC for a given peptide sequence and bacteria using the APEX 1.1 model proposed by Wan et al. [66]. To frame the problem as maximization, we optimize $\text{MIC}$. We seek $K=4$ peptides that together form a potent set of antibiotics for all $T=11$ bacteria. To enable optimization over peptides, we use the VAE model pre-trained on $4.5$ million amino acid sequences from Torres et al. [60] to map the peptide sequence search space to a continuous $256$ dimensional space.
@@
We evaluate MOCOBO on two variations of the peptide design task: “template free" (TF) and “template constrained" (TC). For TF, we allow the optimizer to propose any sequence of amino acids. For TC, we add a constraint that any sequence proposed by the optimizer must have a minimum of $75$ percent sequence similarity to at least one of the $10$ template amino acid sequences in Table˜A.8. These $10$ templates were mined from extinct organisms and selected by Wan et al. [66]. The motivation of the template constrained task is to design peptides specifically likely to evade antibiotic resistance by producing “extinct-like” peptides that bacteria have not encountered in nature in thousands of years. For MOCOBO and all BO baselines (TuRBO, LOL-BO, ROBOT, and MORBO), we handle the optimization constraint by adapting techniques from SCBO [16]. For CluSO and random sampling, we handle the constraint with rejection sampling.
@@
Figure 3: (Left) In vitro results for the two best “template free" (TF1, TF2) and two best “template constrained" (TC1, TC2) runs of MOCOBO for the peptide design task. Columns are the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target pathogenic bacteria $\ldots, B1$1 listed in Table˜A.7. (-) and (+) indicate Gram negative and Gram positive respectively. TF1 and TC1 correspond to the single runs of MOCOBO shown in Table˜1 and Table˜A.1 respectively. Methods used to obtain in vitro MICs are provided in Section˜A.1. (Right) Ablation study comparing MOCOBO to optimization performance where a known “good” partitioning of the $T$ objectives into $K$ subsets is available in advance. We individually optimize $K$ solutions, one for each partition.
@@
In high dynamic range (HDR) images, some pixels (often associated with light sources) can dominate overall contrast, requiring adjustments to reveal detail in low-contrast areas, a problem known as tone mapping [51, Section 6]. Tone mapping algorithms involve various tunable parameters, resulting in a high-dimensional optimization problem of subjectively perceived quality. We seek a covering set of $K=4$ solutions to optimize a set of $T=7$ image aesthetic (IAA) and quality (IQA) assessment metrics from the pyiqa library [9] (see metrics listed in Table˜A.5). Our practical goal is that, while we do not know a priori which metric is best for a particular image, covering all metrics may result in at least one high quality image. We optimize over the $13$-dimensional parameter space of an established tone mapping pipeline to tone-map the “Stanford Memorial Church" [12] and “desk lamp" [7] benchmark images. See Section˜A.8.1 for details.
@@ 4.2 Optimization results
In Figure˜2, we provide optimization results comparing MOCOBO to the baselines discussed above on all tasks. The results show that MOCOBO finds sets of $K$ solutions that achieve higher coverage scores across tasks. The “T Individually Optimized Solutions" baseline appears as a horizontal dotted line in all plots of Figure˜2, representing the average coverage score of $T$ individually optimized solutions, serving as an approximation of the best possible performance without the constraint of a limited $K<T$ solution set. Results in Figure˜2 demonstrate that the smaller set of $K<T$ solutions identified by MOCOBO nearly equals the performance of the complete set of $T$ individually optimized solutions. This result depends on using domain knowledge to choose $K$ large enough to achieve it. Results in Figure˜2 show that matching the performance of $T$ optimized solutions is possible with some values of $K\ll T$.
@@
In Table˜1, we provide the $K=4$ peptides found by one run of MOCOBO for the template free (TF) peptide design task. For each peptide, we provide the the APEX 1.1 model’s predicted MIC for each of the $11$ target bacteria. MIC values $\leq 16$ $\mu$ mol $L^{-1}$ are considered to be “highly active" against the target pathogenic bacteria [66]. We highlight in Table˜1 the comparison between the second peptide with the other three peptides. B1-B7 are Gram negative (GN) bacteria, while B8-B11 are Gram positive (GP). Peptide 2 specialized to the GP bacteria (B8-B11 scores predicted highly active compared to the other peptides) at the expense of broad spectrum activity for the GN bacteria (B5, B6, B7 predicted inactive). By specializing its solutions to GN and GP bacteria separately, MOCOBO achieves low MIC across the target bacteria with only $K=4$ peptides. A similar table of results is also provided for the template constrained (TC) peptide design task (see Table˜A.1).
@@
In Figure˜3 left, we provide in vitro results for the two best TF and two best TC runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. For each run, Figure˜3 left provides the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target bacteria. Results demonstrate that solutions found by MOCOBO optimizing against the in-silico APEX 1.1 model achieve good coverage of the $11$ target pathogenic bacteria in vitro. In Figure˜A.1, we provide the full set of in vitro results for each these runs of MOCOBO, with MIC values obtained by each of the $K=4$ peptides found for each target bacteria. Methods used to obtain in vitro MICs are provided in Section˜A.1.
@@
In Table˜A.2, we provide results for the $K=3$ molecules found by one run of MOCOBO for the molecule design task. Each of the $6$ target elements is successfully present in one of the $K=3$ molecules in the best covering set. These three molecules therefore effectively cover the $T=6$ objectives, with each objective having a max score >0.9>0.9.
@@
In Figure˜4 left, we depict the $4$ varied obstacle courses used for the $T=4$ variation of the rover task and $K=2$ covering trajectories found by a single run of MOCOBO. In Figure˜A.3 and A.4, we provide analogous figures for the $T=8$ and $T=12$ variations. In each example, the MOCOBO optimized set of trajectories covers all obstacle courses such that all obstacles are avoided.
@@
In Figure˜4 right and Figure˜A.2, we provide the original hdr images and the images obtained by the $K=4$ solutions found by a single run of MOCOBO for the church and desk test images respectively. In both variations, three of the four solutions result in high quality images, while one results in a poor quality image (see middle right church image and rightmost desk image). The poor quality images were each generated by the one solution that MOCOBO used to cover metrics $5$ and $6$ (see metric ID numbers in Table˜A.5). This indicates that metrics $5$ and $6$ are poor indicators of true image quality in this setting. This result highlights a useful application of MOCOBO: By dedicating one of the $K$ solutions to maximizing the misleading metrics ($5$ and $6$), the other three solutions to can focus on the remaining subjectively better metrics.
@@ 4.3 Ablation study
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (1) reduces to optimizing the sum of objectives in each relevant subset independently. In Figure˜3 right, we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization. See Section˜A.2.2 for an additional ablation study of MOCOBO when all objectives are highly conflicting.
@@ 5 Discussion
By bridging the gap between traditional multi-objective Bayesian optimization (BO) and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions. See Section˜A.4 for additional discussion and limitations of MOCOBO.
@@ A.2.1 Additional covering sets of solutions found by MOCOBO
In this section, we provide additional examples of covering sets of solutions found by MOCOBO for various tasks from Section˜4.
Figure A.1: In vitro results for the two best “template free" (TF1, TF2) and two best “template constrained" (TC1, TC2) runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. Each row is a single peptide found by a run of MOCOBO. Row TFii-$j$ indicates template free run $i$, peptide $j$. Similarly, TCii-$j$ indicates template constrained run $i$, peptide $j$. Columns are the in vitro MICs for each target pathogenic bacteria $\ldots, B1$1 listed in Table˜A.7. (-) and (+) indicate Gram negative and Gram positive respectively. TF1 and TC1 correspond to the single runs of MOCOBO shown in Table˜1 and Table˜A.1 respectively. Methods used to obtain in vitro MICs are provided in Section˜A.1.
In Figure˜A.1, we provide In vitro results for the two best “template free" and two best “template constrained" runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. Figure˜A.1 provides in vitro MICs for each of the $K=4$ peptides found by each of these runs of MOCOBO, for each of the $11$ target pathogenic bacteria. Methods used to obtain in vitro MICs are provided in Section˜A.1.
@@
Table A.1: The best set of $K=4$ peptide sequences found by one run of MOCOBO for the “template free" peptide design task described in Section˜4.1. For each of the four sequences, we provide the MIC according to the APEX model for each of the $11$ target pathogenic bacteria. The target pathogenic bacteria $\ldots, B1$1 are listed in Table˜A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TC1" in Figure˜3 left, and Figure˜A.1 for in vitro MICs for this set of $K=4$ peptides.
In Table˜A.1, we provide an example of a covering set of peptides found by MOCOBO for the “template constrained" variation of the peptide design task.
In Figure˜A.2, we provide the original hdr image (leftmost panel), and the $K=4$ images produced using the $K=4$ solutions found by a single run of MOCOBO for the church image variation of the image tone mapping task. An analogous result for the desk image variation of the image tone mapping task can be found in the main text in Figure˜4 right. A notable apparent limitation of the existing IAA/IQA metrics we used for the image tone mapping tasks is that they all favored monochromatic images for both the church and desk images. As such, the tone-mapped images obtained appear less colorful than their hand-tuned counterparts reported in prior work.
In Table˜A.2, we provide an example of a covering set of $K=3$ molecules found by MOCOBO for the molecule design task. As mentioned in Section˜4, these three molecules effectively cover the $T=6$ objectives, as evidenced by the presence of all $T=6$ target elements in one of the $K=3$ molecules designed by MOCOBO.
In Figure˜A.3 and Figure˜A.4, we provide examples of a covering set of trajectories found for the $T=8$ and $T=12$ variations of the rover task respectively. In each figure, a panel is shown for each of the $T$ obstacles courses, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. Each panel also shows the best among the $K$ trajectories found by a run of MOCOBO for navigating each obstacle course. An analogous plot for the $T=4$ variation of the rover optimization task is provided in the main text in Figure˜4.
@@
Table A.2: The best set of $K=3$ molecules found by one run of MOCOBO for the Ranolazine MPO multiple element molecule design task described in Section˜4.1. For each of the three molecules, we provide the objective value obtained for each of the $T=6$ objectives. The best/highest objective value is in bold in each column. Each objective aims to add a different target element to Ranolazine (F, Cl, Br, Se, S, and P). The $T=6$ target elements are in bold in the smiles string [69] representation of each molecule.
@@ A.2.2 Additional ablation study: $T$ highly conflicting objectives
MOCOBO is designed for cases where some, but not all, of the objectives are highly conflicting, and this is the case for all of the tasks considered in Section˜4. The conflicting pairs of the objectives prevent a single solution from optimizing all objectives well. The fact that not all pairs of objectives are completely conflicting, is the reason why it is possible to cover all T objectives with a small set of K solutions. If all pairs of the T objectives are highly conflicting, there is by definition no possible set of K<T solutions such that all objectives are well optimized. In this case, it’s best to use T individually optimized solutions if your goal is to find at least one solution that well optimizes each objective. However, it is still interesting to consider the performance of MOCOBO in the setting of T objectives that are pairwise highly conflicting. To investigate this, we construct a new variation of the multi-objective rover task described in Section˜4.1. We design T=3 obstacle courses such that it is impossible for the rover to take any single path that avoids all obstacles in any pair of the obstacle courses. This results in T=3 objectives that are pairwise highly conflicting. We then run 20 replications of MOCOBO on this problem with K=2, asking MOCOBO to design K=2 solutions that cover the T=3 pairwise highly conflicting objectives. For this task, MOCOBO got an average coverage score of $2.932$, while T individually optimized solutions got an average coverage score of $12.899$. It is unsurprising that T individually optimized solutions achieved better coverage here since all three objectives are pairwise highly conflicting. However, this result demonstrates that MOCOBO is able to design sets of K=2 solutions that still achieve fairly high coverage scores.
@@
Figure˜A.5 provides a diagram of the T=3 pairwise highly conflicting obstacle courses, and an example of one of the covering sets of K=2 trajectories found by a single run of MOCOBO. One optimized trajectory (shown in blue) is designed by MOCOBO such that it specializes to navigate the first obstacle course, navigating it without hitting any obstacles (Leftmost Panel). The other optimized trajectory (shown in magenta) is designed by MOCOBO such that it avoids all obstacles in the third obstacle course (Rightmost Panel), while also minimizing total amount of impact with the obstacles in the second obstacle course (Middle Panel). Since it is by-design impossible to avoid all obstacles in all T=3 obstacle courses with only K=2 solutions, Figure˜A.5 demonstrates that MOCOBO was able to successfully balance trade-offs among the objectives, designing a set of K=2 solutions that minimized the total amount of obstacle impact across the three obstacle courses.
A.2.3 Wall-clock runtime of Algorithm˜1
In Table˜A.3, we provide an empirical evaluation of the average wall-clock runtime of Algorithm˜1 with different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). We provide average run times with $N=20,200,2000,20000,200000,$ and $2000000$ for each combination of $K$ and $T$ that we ran MOCOBO on in Section˜4. To improve the speed of Algorithm˜1, in all experiments, we compute the inner for loop (lines 5-6 in Algorithm˜1) in parallel. This involves simply computing the marginal coverage of adding all $N$ data points at once in parallel, and then updating $A$ with the point that achieved maximum marginal converge improvement. Parallel computation of the marginal coverage improvement of the $N$ points is straightforward to implement using PyTorch. In Table˜A.3, the “Runtime" column gives the average runtime of Algorithm˜1 with this parallel computation of the of the inner for loop, and the “Runtime No Parallelization" column gives the average runtime without the parallel computation. Notice that as $N$ grows large, this parallel computation becomes essential to achieve the reasonably fast run times of Algorithm˜1 needed to run MOCOBO. Average run times reported in Table˜A.3 are computed by running Algorithm˜1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs.